OpenStudy (anonymous):

the fourth term of the sequence is 19 and the eleventh term is 54. find the eighth term.

OpenStudy (anonymous):

look like an arithmetic series: common difference = 54-19 / 11-4 = 35/7 = 5 eight term = 19 + 4(5) = 39

OpenStudy (king):

\[a _{n}=a _{1}+(n-1)d\]

OpenStudy (king):

so \[a _{4}=a _{1}+3d=19\]

OpenStudy (king):

and \[a _{3}=a _{1}+2d=54\]

OpenStudy (king):

so a1=54-2d=19-3d a1=-35 d=18

OpenStudy (king):

so a11=-35+10*18 =-35+180 =145

OpenStudy (anonymous):

am not sure u got it .......pls check the question again